Find the dimensions of a rectangle with an area of sixty square feet and a diagonal of thirteen feet.
Found 2 solutions by edjones, jerryguo41:
Answer by edjones(8007)

Let the length=a and the width=b
ab=60
a=60/b
a^2+b^2=13^2
(60/b)^2+b^2=169 Substitute 60/b for a
3600/b^2 + b^2 = 169
3600+b^4=169b^2 multiply each side by b^2
b^4-169b^2+3600=0
(b^2-144)(b^2-25)=0
(b+12)(b-12)(b+5)(b-5)=0
b=5', a=12'
.
Ed

Answer by jerryguo41(197)